# Cost Efficient Fairness Audit Under Partial Feedback

### v1

@@ 1 Introduction
Machine learning models can amplify and exacerbate biases present in their training data. To address this, fairness audits and interventions have emerged as key topics in the fair machine learning literature MMS+ (21); BHN (23); YZC+ (24). Consequently, a wide range of fairness metrics DHP+ (12); HPS (16); GVF (20), fairness interventions BDH+ (19); BDE+ (20), and audit frameworks have been proposed and adopted SKH+ (18); HWVDI+ (19).
@@
This work addresses the problem of auditing machine learning models for fairness, an issue of growing significance in decision-making systems BG (18); LB (23); MSKF (24). Previous work on fairness audit covers a range of individual and group fairness metrics XYS (20); KNRW (18) and construction of statistical tests for auditing TBKN (21); SMBN (21). In the simplest fairness audit setting, a decision-maker employs a classifier $f$ that takes as input a feature vector $x$ and a sensitive attribute or group label $a$ and makes a binary prediction $f(x,a)$. An auditor aims to estimate the fairness of $f$ using a standard fairness metric like Equalized Odds, Equal Opportunity with respect to sensitive attributes.
@@
In many real-world settings such as a bank disbursing loans using a classifier $f$, the class label $y$ indicating the loan repayment status is available only for applicants who are offered loans (i.e., when $f(x,a)=1$). Thus, the historical data is biased or incomplete, whereas any new data exploration can be costly (e.g., random disbursal of loans has higher defaults). Previous work under this incomplete data model can be broadly categorized into two settings. First, learning fair classifiers from incomplete labeled data (along with unlabeled data) without actively acquiring the full set of labels LKL+ (17); KZ (18); KRS+ (20); Wei (21). Second, online fair classification with partial feedback, where labels are revealed only for positively classified data points BLR+ (19); KMC (24). Complementary to these works, we study the fairness audit problem in the partial feedback setting under a novel cost model that incorporates both access to historically biased or incomplete data and the cost of acquiring additional labels to complete it.
@@
In the partial feedback setting, due to the unavailability of labels for negatively classified individuals ($f(x,a)=0$), any fairness audit algorithm must acquire these labels to complete the data and accurately estimate outcome-based fairness metrics such as equal opportunity and equalized odds on the true data distribution. Notably, in scenarios such as bank loan applications, granting a loan to a previously negatively classified individual may incur a cost if the individual fails to repay. This cost is not modeled in prior fairness audit work, which primarily focuses on hypothesis testing for a given unfairness tolerance KNRW (18); XYS (20).
Our Contributions: We study the problem of fairness audit under partial feedback with a novel cost model. Our cost model captures a central concern of real-world scenarios, where there is a price to pay when the auditor decides to approve an applicant with negative outcome (e.g., loan default). For our cost model, we provide a complete characterization of the sample and cost complexity required for auditing fairness in the non-parametric setting by establishing a lower bound and proposing an audit algorithm that matches this bound up to logarithmic factors.
@@
To further investigate auditing with partial feedback, we study a structured setting where the group-wise distribution of features and labels follows a mixture of exponential families, a common analytical assumption in the fairness literature DWY+ (20); MWL (22). In this setting, we provide a novel perspective by viewing the classifier $f$ as inducing a truncation on the input distribution by positively classifying only a subset of samples. We leverage tools from learning with truncated samples DGTZ (18); LWZ (23) to address the fairness audit problem under partial feedback. Furthermore, we generalize prior work on maximum a-posteriori (MAP) oracles for mixtures of spherical Gaussians Mah (11) to mixtures of exponential families, which may be of independent interest.
@@ 1.1 Related work
Learning with Partial Feedback: The problem of learning from selection-biased or partial feedback data has been rigorously studied in various contexts in past works. LKL+ (17) study the problem of evaluation with selection bias and propose using the heterogeneity of available human decision makers to account for unobserved data. KZ (18) study the residual unfairness of models trained on selection-biased data and propose to correct for sample selection bias by reweighting using benchmark data. DADC (18) leverages the observation that human experts can exhibit consistent prediction in certain regions and uses this to design a data augmentation scheme. KRS+ (20) propose using exploration-based stochastic decision rules over predictive models trained on selection-biased data, to achieve fairness and utility with sample selection bias. Wei (21) study utility maximization under selection bias in an MDP setup. Finally, YLN (22) proposes an active debiasing model that adjusts a classifier initially learned on a partial feedback dataset, and where the predictor evolves over time with bounded exploration.
BLR+ (19); BJW (20); Bec (24) study fair online classification for group and individual fairness metrics under the partial feedback setting. Similarly, KMC (24) studies the problem of data collection and fair classification across all subgroups in a partial feedback setting. Complementary to these learning problems, our objective is to construct a cost-efficient algorithm to audit a decision maker for fairness using a framework that combines past, selection-biased data, and online exploration. To the best of our knowledge, our work is the first to define and minimize cost and sample complexity while testing for equalized odds under partial feedback. We also note our work is different from the work on fairness with missing data ZL (21); MQH+ (25) where the objective is to work with missing covariates but full outcome information, unlike partial feedback.
Fairness Audit: Fairness auditing is a well-studied research topic, with practical toolkits for auditing BDH+ (19); SKH+ (18), quantifying sample sizes SC (23); SXK+ (23), and theoretical frameworks. CCGWR (23) propose a sequential hypothesis testing framework with the ability to ingest data continuously and handle drifts in the data. SMBN (21); TBKN (21); GMD+ (25) leverage optimal transport techniques to construct statistical tests and then enforce exact and approximate fairness. YZ (22) study the auditing problem in a realizable setting via active queries. BSO+ (24) study fairness audit with missing confounders for treatment rate estimation. Other seminal works in fairness auditing literature handle potentially infinite and a rich collection of subgroups KNRW (18); MOW+ (19); CC (24). To the best of our knowledge, we are the first to study fairness audit under partial feedback.
@@ 2 Problem Formulation
In the fairness audit problem, we are interested in testing the fairness of a given classifier $f:{\mathcal{X}}\times{\mathcal{A}}\to\{0,1\}$. Here $\mathcal{X}$} denotes the set of features and $\mathcal{A}$} denotes the group labels of individuals. Every individual is characterized by a random tuple $(X,A,Y)$, $X\in{\mathcal{X}}$, $A\in{\mathcal{A}}$ and $Y\in\{0,1\}$, where $Y$ is the true label of the individual. If $f(X,A)=1$ for an individual, the individual is classified as positive; otherwise, they are classified as negative. The distribution $\mathcal{D}$} of $(X,A,Y)$ is unknown to the auditor. We adopt equalized odds as our primary fairness notion for the remainder of the paper, although our techniques can also be applied to other fairness metrics such as demographic parity DHP+ (12) and equal opportunity HPS (16).
Throughout the paper, we use the shorthand notation $f\coloneqq f(X,A)$ whenever there is no ambiguity regarding $X$ and $A$. We define $\mathbb{P}[f\mid Y,A]\coloneqq\mathbb{P}_{X}[f(X,A)\mid Y,A].$. The equalized odds difference (EOD) is then defined as  HPS (16):
@@
Under this notion, a classifier $f$ is said to be fair if $\Delta=0$ and unfair if $\Delta>\varepsilon$ for some $\varepsilon>0$. We pose the fairness audit of $f$ as a hypothesis testing problem that we now describe.
Definition 1 ($(\varepsilon,\delta)$-Fairness Audit).
@@
Given $\varepsilon,\delta\in(0,1)$, an algorithm/auditor is said to succeed in the $(\varepsilon,\delta)$-fairness audit if with probability at least $1-\delta$ it can identify correctly which of the following hypotheses is true: (i) $\mathsf{FAIR}:\Delta=0$, OR (ii) $\mathsf{UNFAIR}:\Delta>\varepsilon$.
@@
The parameter $\varepsilon$ can be seen as a fairness threshold which the classifier must comply with and whose value can be chosen according to the requirement of the problem.
@@
Partial Feedback: We now describe the $(\varepsilon,\delta)$-fairness audit under the partial feedback (PF) model. In the vanilla PF model, the true label is revealed for positively classified individuals only, e.g., the loan default ($Y\in\{0,1\}$) can only be observed after the loan is given ($f=1$) by the bank. However, we allow the auditor to request true labels by overriding the negative classification made by the decision-maker. Hence the auditor can explore additional data, which is necessary for the following reason. In our model, an auditor can access $(X,A,Y)$ of individuals who were classified positively in the past by $f$, but only $(X,A)$ of individuals who were classified negatively. From this database, the probability $\mathbb{P}[f=1,Y=y,A=a]$, can be estimated, but not $\mathbb{P}[f=0,Y=y,A=a]$, $y\in\{0,1\}$ and $a\in{\mathcal{A}}$, since $Y$ is unknown when $f=0$. We therefore let the auditor access online samples from $\mathcal{D}$} and request $X$ and $Y$ of negatively classified individuals. With this combination of past and online data, the auditor must succeed in the $(\varepsilon,\delta)$-fairness audit.
@@
Cost model: We now introduce a novel but natural cost model motivated by practical examples of the partial feedback setting like bank loan applications KHD (22), criminal recidivism ALMK (22), and college admissions CDGN+ (23). The auditor is charged a fixed cost $c_{feat}\geq 0$ every time they request only the feature of an individual for whom $f=0$. If the auditor requests only the true label or both the feature and the true label of such an individual, they are charged $c_{feat}+c_{lab}\cdot\mathds{1}\{Y=0\}$, $c_{lab}\geq 0$. We illustrate the motivation of this cost definition through the example of bank loan application. Firstly, there may be a cost involved in collecting the data of an individual, which is captured by $c_{feat}$. Secondly, to obtain the true label, the auditor has to force the bank (since $f=0$) to give loan to the individual. If the individual fails to return the loan $(Y=0)$, only then the bank suffers a monetary loss. To capture this asymmetry, we define the cost formally below.
@@
Let the auditor conclude the $(\varepsilon,\delta)$-fairness audit after running their algorithm on $N$ iid individuals sampled from $\mathcal{D}$} and let $g_{i}$ and $h_{i}$ respectively denote whether the auditor observes the feature only and the true label of the $i$-th individual. Then, the audit cost is:
@@
Typically, $c_{feat}\ll c_{lab}$, e.g. in the bank loan application, the cost of loan default is much higher than the loan processing fee, credit risk assessment etc. For the partial feedback setting, the audit cost is a more effective performance measure than the usual sample complexity in capturing the asymmetry in the cost of true label realization. Further, since the auditor might request features and labels separately, it is not clear what should count towards sample complexity. To this end, the cost definition provides a unified measure of performance of fairness audit algorithms in the partial feedback setting. Therefore, with this novel definition of audit cost, we now state the main question that we study in this work:
Design an algorithm that succeeds in the $(\varepsilon,\delta)$-fairness audit with as small audit cost as possible.
@@
In our mixture model, the class of distribution from the exponential family is known (i.e., $h$, $T$ and $W$ are known). The data $(X,A,Y)$ is generated is as follows: first a group label $A$ is sampled, then true label $Y$ is sampled as $Y\sim Ber(q_{1\mid A})$ and finally, feature $X$ is sampled as $\mathbb{P}[X\mid Y,A]={\mathcal{E}}_{\theta^{*}_{Y,A}}(X)=h(X)\exp(\theta^{*\intercal}_{Y,A}T(X)-W(\theta^{*}_{Y,A}))$, where $\theta^{*}_{y,a}\in\Theta\subset\mathbb{R}^{d}$, $y\in\{0,1\}$, $a\in{\mathcal{A}}$, are unknown. Here $\Theta$ is a known convex set.
In the following section, we highlight our main results. First we bound the audit cost of a natural baseline algorithm for the blackbox model. We then state another algorithm for this setting that has even lower audit cost. Moreover, we also show that this algorithm is tight in terms of the number of true labels it requests by proving a lower bound. Finally, we give another novel algorithm for the mixture model that achieves significantly lower audit cost than the blackbox case. This algorithm has two main novelties. Firstly, it views the past database as truncated samples from the true distribution thus leveraging techniques from learning from truncated samples literature DGTZ (18); LWZ (23). Secondly, it generalizes existing work on learning mixture of distributions from spherical Gaussians Mah (11) to general exponential family.

### v2

@@ 1 Introduction
Machine learning models can amplify and exacerbate biases present in their training data. To address this, fairness audits and interventions have emerged as key topics in the fair machine learning literature Mehrabi et al. (2021); Barocas et al. (2023); Ye et al. (2024). Consequently, different fairness metrics Dwork et al. (2012); Hardt et al. (2016); Garg et al. (2020), fairness interventions Bellamy et al. (2019); Bird et al. (2020), and audit frameworks have been proposed and adopted Saleiro et al. (2018); Holstein et al. (2019).
@@
This work addresses the problem of auditing machine learning models for fairness, an issue of growing significance in decision-making systems Buolamwini and Gebru (2018); Landers and Behrend (2023); Mökander et al. (2024). Previous work on fairness audit covers a range of individual and group fairness metrics Xue et al. (2020); Kearns et al. (2018) and construction of statistical tests for auditing Taskesen et al. (2021); Si et al. (2021). In the simplest fairness audit setting, a decision-maker employs a classifier $f$ that takes as input a feature vector $x$ and a sensitive attribute or group label $a$ and makes a binary prediction $f(x,a)$. An auditor aims to estimate the fairness of $f$ for a standard fairness metric such as Equal Opportunity (i.e., equal True Positive Rates for different groups) using the available data.
@@
As a motivating real-world example, consider loan disbursement using a classifier $f$ for approval, where the class label $y$ indicating the loan repayment status is available only for applicants who are offered loans (i.e., when $f(x,a)=1$). Thus, the historical data is biased or incomplete, whereas any new data exploration can be costly (e.g., random disbursal may have higher defaults). Previous work under this incomplete data model can be broadly categorized into two settings. First, semi-supervised fair learning from incomplete labeled data and complete unlabeled data, without actively acquiring additional labels Lakkaraju et al. (2017); Kallus and Zhou (2018); Kilbertus et al. (2020); Wei (2021). Second, online fair classification with partial feedback, where labels are revealed only for positively classified data points Bechavod et al. (2019); Keswani et al. (2024). Complementary to these works, we study the fairness audit problem in the partial feedback setting under a novel cost model that incorporates both the access to historically biased or incomplete labeled data and the cost of acquiring additional labels on complete but unlabeled data.
@@
In the partial feedback setting, due to unavailability of labels for negatively classified individuals ($f(x,a)=0$), any fairness audit algorithm must acquire these labels to complete the data and accurately estimate outcome-based fairness metrics such as equal opportunity and equalized odds on the true data distribution. Notably, in scenarios such as bank loan applications, granting a loan to a previously negatively classified individual may incur a cost if the individual fails to repay. This cost is not modeled in prior fairness audit work primarily focusing on hypothesis testing for a given unfairness tolerance Kearns et al. (2018); Xue et al. (2020).
1.1 Our Contributions
$\bullet$ We initiate the study of fairness audit under partial feedback, an important setting that captures several real-world scenarios. We introduce a novel cost model for data exploration to capture a central concern, namely, the audit or the self-assessment incurs a cost when the labels requested on additional unlabeled data turn out to be negative (e.g., loan default, unfavorable outcomes in randomized control trials).
$\bullet$ For our cost model, we provide a complete characterization of the sample and cost complexity required for auditing fairness in the non-parametric setting by establishing a lower bound (Theorem 3) and proposing an audit algorithm that matches this bound up to logarithmic factors (Theorem 2). The proposed algorithm provably improves over the audit cost of a natural baseline algorithm (Theorem 1, Remark 1).
$\bullet$ To further investigate auditing with partial feedback, we study a structured setting where the group-wise distribution of features and labels follows a mixture of exponential families, a common analytical assumption in the fair learning Dutta et al. (2020); Ma et al. (2022). In this setting, we design a new algorithm that achieves significantly lower audit cost under mild assumptions (Theorem 4).
$\bullet$ Our algorithms and their cost guarantees are agnostic to the hypothesis class and work for any classifier whose fairness is to be audited. We empirically study them on standard real-world datasets in fair classification (e.g., Adult Income, Law School) for logistic regression classifiers with significantly improved cost of audit or self-assessment.
1.2 Important Remarks on Our Contributions
$\bullet$ To the best of our knowledge, this work is the first to address the important problem of fairness audit under the partial feedback setting capturing several real-world scenarios. While the problems of learning fair classifiers in partial feedback, and of fairness audit with full data have been studied (see next section), through this work, we aim to provide a principled framework for fairness audit that accounts for the cost of additional data in the partial feedback setting.
$\bullet$ Our results in the blackbox setting require several novel techniques both in terms of algorithm design and analysis. To establish the cost lower bound (Theorem 3), we formalize the sequential interaction of the auditor with the online partial feedback data and derive a KL-divergence decomposition lemma (Lemma 7, Appendix A.3). Furthermore, we carefully design two (family of) instances that are close in KL-divergence but only one of them is fair to conclude the lower bound. For the design of algorithm, we rely on the design of a stopping time and analyse it via a novel use of the concentration of the Negative Binomial distribution instead of the standard Chernoff-Hoeffding type bounds.
@@ 1.2 Important Remarks on Our Contributions
$\bullet$ In the mixture model setting, we provide a novel perspective by viewing the classifier $f$ as inducing a truncation on the input distribution by positively classifying only a subset of samples. This viewpoint allows us to leverage tools from learning with truncated samples Daskalakis et al. (2018); Lee et al. (2023) to address the fairness audit problem under partial feedback. We believe this novel connection can be of wider interest, for example in understanding the truncation arising from popular fair classification methods. Furthermore, we generalize prior work on maximum a-posteriori (MAP) oracles for mixtures of spherical Gaussians Mahalanabis (2011) to handle mixtures of exponential families, which may be of independent interest.
@@ 1.3 Related work
Learning with Partial Feedback: The problem of learning from selection-biased or partial feedback data has been rigorously studied in various contexts in past works. Lakkaraju et al. (2017) study the problem of evaluation with selection bias and propose using the heterogeneity of available human decision makers to account for unobserved data. Kallus and Zhou (2018) study the residual unfairness of models trained on selection-biased data and propose to correct for sample selection bias by reweighting using benchmark data. De-Arteaga et al. (2018) leverages the observation that human experts can exhibit consistent prediction in certain regions and uses this to design a data augmentation scheme. Kilbertus et al. (2020) propose using exploration-based stochastic decision rules over predictive models trained on selection-biased data, to achieve fairness and utility with sample selection bias. Wei (2021) study utility maximization under selection bias in an MDP setup. Finally, Yang et al. (2022) proposes an active debiasing model that adjusts a classifier initially learned on a partial feedback data, with the predictor evolving over time with bounded exploration.
Bechavod et al. (2019, 2020); Bechavod (2024) study fair online classification for group and individual fairness metrics under the partial feedback setting. Similarly, Keswani et al. (2024) studies the problem of data collection and fair classification across all subgroups in a partial feedback setting. Complementary to these learning problems, our objective is to construct a cost-efficient algorithm to audit a decision maker for fairness using a framework that combines past, selection-biased data, and online exploration. To the best of our knowledge, our work is the first to define and minimize cost and sample complexity while testing for equalized odds under partial feedback. We also note our work is different from the work on fairness with missing data Zhang and Long (2021); Miao et al. (2025) where the objective is to work with missing covariates but full outcome information, unlike partial feedback.
Fairness Audit: Fairness auditing is a well-studied research area, with practical toolkits for auditing Bellamy et al. (2019); Saleiro et al. (2018), quantifying sample sizes Singh and Chunara (2023); Singh et al. (2023), and theoretical frameworks. Chugg et al. (2023) propose a sequential hypothesis testing framework with the ability to ingest data continuously and handle drifts in the data. Si et al. (2021); Taskesen et al. (2021); Ghassemi et al. (2025) leverage optimal transport techniques to construct statistical tests and then enforce exact and approximate fairness. Yan and Zhang (2022) study auditing in a realizable setting via active queries. Byun et al. (2024) study fairness audit with missing confounders for treatment rate estimation. Other seminal works in fairness auditing handle potentially infinite and a rich collection of subgroups Kearns et al. (2018); Morina et al. (2019); Cherian and Candès (2024). To the best of our knowledge, we are the first to study fairness audit under partial feedback.
@@ 2 Problem Formulation
In the fairness audit problem, we are interested in testing the fairness of a given classifier $f:{\mathcal{X}}\times{\mathcal{A}}\to\{0,1\}$. Here $\mathcal{X}$} denotes the set of features and $\mathcal{A}$} denotes the group labels of individuals. Every individual is characterized by a random tuple $(X,A,Y)$, $X\in{\mathcal{X}}$, $A\in{\mathcal{A}}$ and $Y\in\{0,1\}$, where $Y$ is the true label of the individual. If $f(X,A)=1$ for an individual, the individual is classified as positive; otherwise, they are classified as negative. The distribution $\mathcal{D}$} of $(X,A,Y)$ is unknown to the auditor. We adopt equalized odds as our primary fairness notion for the remainder of the paper, although our techniques can also be applied to other fairness metrics such as demographic parity Dwork et al. (2012) and equal opportunity Hardt et al. (2016).
Throughout the paper, we use the shorthand notation $f\coloneqq f(X,A)$ whenever there is no ambiguity regarding $X$ and $A$. We define $\mathbb{P}[f\mid Y,A]\coloneqq\mathbb{P}_{X}[f(X,A)\mid Y,A]$. The equalized odds difference is defined as  Hardt et al. (2016):
@@
Under this notion, a classifier $f$ is said to be fair if $\Delta\leq\gamma$ and unfair if $\Delta>\gamma+\varepsilon$ for some $\gamma\geq 0,\varepsilon>0$. We pose the fairness audit of $f$ as a hypothesis testing problem:
Definition 1 ($(\gamma,\varepsilon,\delta)$-Fairness Audit).
@@
Given $\gamma\geq 0$, $\varepsilon,\delta\in(0,1)$, an algorithm/auditor is said to succeed in the $(\varepsilon,\delta)$-fairness audit if with probability at least $1-\delta$ it can identify correctly which of the following hypotheses is true: (i) $\mathsf{FAIR}:\Delta\leq\gamma$, OR (ii) $\mathsf{UNFAIR}:\Delta>\gamma+\varepsilon$.
@@
The parameter $\gamma$ can be seen as a fairness threshold which the classifier must comply with and whose value can be chosen according to the requirement of the problem.
@@
Partial Feedback: We now describe the $(\gamma,\varepsilon,\delta)$-fairness audit under the partial feedback (PF) model. In the vanilla PF model, the true label is revealed for positively classified individuals only, e.g., the loan default ($Y\in\{0,1\}$) can only be observed after the loan is given ($f=1$) by the bank. However, we allow the auditor to request true labels by overriding the negative classification made by the decision-maker. Hence the auditor can explore additional data, which is necessary for the following reason. In our model, an auditor can access $(X,A,Y)$ of individuals who were classified positively in the past by $f$, but only $(X,A)$ of individuals who were classified negatively. From this database, the probability $\mathbb{P}[f=1,Y=y,A=a]$, can be estimated, but not $\mathbb{P}[f=0,Y=y,A=a]$, since $Y$ is unknown when $f=0$. Hence, we let the auditor access online samples from $\mathcal{D}$} and request $X$ and $Y$ of negatively classified individuals. With this blend of past and online data, the auditor must succeed in $(\gamma,\varepsilon,\delta)$-fairness audit.
@@
Cost model: We now introduce a novel but natural cost model for exploring online data motivated by practical examples of the partial feedback setting like bank loan applications Kumar et al. (2022), criminal recidivism Angwin et al. (2022), and college admissions Corbett-Davies et al. (2023). The auditor can observe the sensitive attribute of the individual but is charged a fixed cost $c_{feat}\geq 0$ every time they request only the feature of an individual for whom $f=0$. If the auditor requests only the true label or both the feature and the true label of such an individual, they are charged $c_{feat}+c_{lab}\cdot\mathds{1}\{Y=0\}$, $c_{lab}\geq 0$. We illustrate the motivation of this cost definition through the example of bank loan application. Firstly, there may be a cost involved in collecting the personal details of an individual, which is captured by $c_{feat}$ and incurred to the auditor when the bank does not approve the loan. Secondly, to obtain the true label, the auditor has to force the bank (since $f=0$) to give loan to the individual. If the individual fails to return the loan $(Y=0)$, only then the bank suffers a monetary loss. To capture this asymmetry, we define the cost formally below.
@@
Let the auditor conclude the $(\gamma,\varepsilon,\delta)$-fairness audit after running their algorithm on $N$ iid individuals sampled from $\mathcal{D}$} and let $g_{i}$ and $h_{i}$ respectively denote whether the auditor observes the feature only, and the true label of the $i$-th individual. Then, the audit cost is:
@@
Typically, $c_{feat}\ll c_{lab}$, e.g. in the bank loan application, the cost of loan default is much higher than the loan processing fee, credit risk assessment etc. For the partial feedback setting, the audit cost is a more effective performance measure than the usual sample complexity in capturing the asymmetry in the cost of true label realization. Further, since the auditor might request features and labels separately, it is not clear what should count towards sample complexity. To this end, the cost definition provides a unified measure of performance of fairness audit algorithms in the partial feedback setting. Hence, with this novel definition of audit cost, we now state the main question studied in this work:
Design an algorithm that succeeds in the $(\gamma,\varepsilon,\delta)$-fairness audit with as small audit cost as possible.
@@
In our mixture model, the class of distribution from the exponential family is known (i.e., $h$, $T$ and $W$ are known). The data $(X,A,Y)$ is generated is as follows: first a group label $A$ is sampled, then true label $Y$ is sampled as $Y\sim\mathrm{Ber}(q_{1\mid A})$ and finally, feature $X$ is sampled as $\mathbb{P}[X\mid Y,A]={\mathcal{E}}_{\theta^{*}_{Y,A}}(X)=h(X)\exp(\theta^{*\intercal}_{Y,A}T(X)-W(\theta^{*}_{Y,A}))$, where $\theta^{*}_{y,a}\in\Theta\subset\mathbb{R}^{d}$, $y\in\{0,1\}$, $a\in{\mathcal{A}}$, are unknown. Here $\Theta$ is a known convex set.
In the following section, we highlight our main results which includes the audit cost of a natural baseline algorithm, followed by an improved cost algorithm for the blackbox model and a matching lower bound. Next, for the mixture model, we design another algorithm that further improves the cost.